OpenStudy (anonymous):

What equation has a slope of -5 and includes the point (-1, -2). A. y+2=3(x-1) B. y-1=3(x+2) C. y+2=-5(x+1) D. none

jigglypuff314 (jigglypuff314):

point slope form: y - y1 = m(x - x1) where m = slope so which equation has slope = -5 = m ?

jigglypuff314 (jigglypuff314):

Please try to avoid giving such direct answers, it is against OpenStudy's code of conduct http://openstudy.com/code-of-conduct This site has a policy for "Give Help, Not Answers" violations of such policies could potentially result in an account suspension or ban. So, please keep that in mind from now on :)

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

okay I will

OpenStudy (anonymous):

^u don't need to worry that much. simply substitute the given coordinates to the equations and see weather the equation holds